The Jammin Company is financed entirely with equity. The company is considering a loan of $3.6 million. The loan will be repaid in equal principal installments over the next two years and has an interest rate of 6 percent. The company’s tax rate is 25 percent.

  

According to MM Proposition I with taxes, what would be the increase in the value of the company after the loan?
Answer #1

Rate per month =6%/12 =0.5%
Number of Periods =12*2 =24
FV of the loan =3600000*(1+0.5%)^24 =4057775.1943

The installments for 24 months =4057775.1943/12=338147.9329
The principal part =3600000/12 =300000
The interest part =338147.93-300000 =38147.9329

The Interest tax shield =Interest*Tax rate =38147.9329*25% =9536.9832

Increase in the value of company after the loan =9536.9832*((1-(1+0.5%)^-24)/0.5%) =215181.68